solve by substitution method (linear equation) x-5y=4 and 3x-4y=12

Answer by AAfter.Com(18)

x-5y=4 => x = 5y + 4
3x-4y=12 => 3(5y + 4) - 4y = 12
=> 15y + 12 - 4y = 12
=> 11y = 0
=> y = 0
Hence, x = 5y + 4 = 5X0 + 4 = 4
Hence, the solution set is (4,0)